a man takes brick out of fire at 285C he drops it in a bucket of water and hen covers the bucket. the bucket and cover are amade of insulating materials. the bucket contains 2.6kg of ice originally at -5C. the 1.3kg brick is made out of iro with a specific heat capacity of 448J/kg.C. find the temp of the brick and water when equilibrium is reached.
1
Expert's answer
2020-02-07T10:33:10-0500

From the conservation of energy:


"Q_1=Mc_i(T_1-T_0)"

"Q_1=2.6(2100(0-(-5)))=27300\ J"

"Q=mc_b(T_b-T_1)=(1.3)(448)(285-0)=165984\ J"

"Q_2=ML=(2.6)(340000)=884000\ J"

"Q<Q_1+Q_2"

Thus, the final temperature of the system will be


"T=0\degree C"